Question: Zahra bought 25 daily news papers and 5 Friday issues in a month and paid rupees 80 for them. Next month she paid rupees 78 for 27 dailies and 3 Friday issues. Find the price of each daily newspaper & each Friday issue.
Answer by Mathtut(3670)

let d and f be the cost of daily and friday newpapers
:
25d+5f=80........eq 1
27d+3f=78........eq 2
:
multiply eq 1 by -3 and eq 2 by 5 and add the two revised equation together
:
-75d-15f=-240......eq 1 revised
135d+15f=390.......eq 2 revised
:
f terms are eliminated and we are left with 135d-75d=390-240
:
60d=150
:
d=2.5 ruppees per daily
:
take d's found value and plug it back into any equation.I choose eq 1
:
25(2.5)+5f=80
:
62.5+5f=80
:
5f=17.5
:
f=3.5 ruppees per friday issue